# Formation of millisecond pulsar-helium star binaries

### v1

@@ 1 Introduction
The recent discovery of the pulsar J1928+1815 by the Five-hundred-meter Aperture Spherical radio Telescope (FAST) offers a novel perspective on the formation and evolution of millisecond pulsars (MSPs, Yang et al., 2025). This pulsar, with a spin period of 10.55 ms, resides in a binary system with an orbital period of approximately 0.15 day. Its companion is thought to be a helium (He) star with mass $1-1.6\,M_{\odot}$.
@@
Observationally, binary systems containing stripped He stars have been identified in the Magellanic Clouds (Drout et al., 2023). In the Milky Way, X-ray binaries with He star donors have also been discovered, with accretors including white dwarfs (WDs, Mereghetti et al., 2009) and neutron stars (NSs, Mereghetti et al., 2016). However, prior to this discovery, no binary systems containing recycled pulsars and He star companions had been identified. The detection of PSR J1928+1815 fills this observational gap, representing the first confirmed binary system of this kind. It provides crucial evidence for testing and extending existing binary evolution theories and opens a new window for in-depth studies of the accretion-induced spin-up process in MSPs. Through detailed observations and theoretical analyses of this system, we can gain a better understanding of the formation mechanisms and physical properties of MSPs.
@@
MSPs are widely believed to form through the recycling of old NSs via accretion of matter and angular momentum from a companion star in low-mass X-ray binaries (LMXBs, Alpar et al., 1982; Radhakrishnan & Srinivasan, 1982). This hypothesis is supported by the detection of millisecond X-ray pulsations in LMXBs and observations of transitions from accreting millisecond X-ray pulsars to MSPs (Wijnands & van der Klis, 1998; Archibald et al., 2009; Papitto et al., 2013; Roy et al., 2015).
@@
In contrast, systems such as PSR J1928+1815 are thought to originate from intermediate-mass or high-mass X-ray binaries (IMXBs or HMXBs) consisting of a NS and a more massive companion star (e.g., Vigna-Gómez et al., 2018; Shao & Li, 2018; Deng et al., 2024). When the companion star evolves to be a giant star and overflows its Roche-lobe, mass transfer (MT) will typically be dynamically unstable because of the large mass ratio. The binary enters the common envelope (CE) phase (Ivanova et al., 2013), during which the giant star’s envelope is shed. If the binary survives CE evolution, a recycled NS and a He star, which is the core of the giant star, are left.
@@
Theoretically, accelerating a NS’s spin period to 10 ms requires accretion of at least $0.01\,M_{\odot}$ of matter (e.g., Tauris et al., 2012), and a NS in intermediate-mass or high-mass X-ray binaries may undergo multiple accretion phases. Initially, there is accretion from the stellar wind of the giant companion before Roche lobe overflow (RLO). While wind accretion can provide some accreted matter ($\sim 0.005\,M_{\odot}$, Yang et al., 2025), its impact on the NS’s spin is uncertain, as it may both accelerate and decelerate the spin (see Tauris & van den Heuvel, 2023; Mao & Li, 2024, and references therein). Current observations of wind-fed X-ray pulsars indicate that the spin periods of NSs range from $\sim 1-10,000$ s (see Weng & Ji, 2024, Figure 1), with a typical value of $\sim 100$ s (Lyne & Graham-Smith, 2012).
@@
Subsequently, there is the RLO phase. During this phase, if the MT rate is constrained by the Eddington limit, accreting $0.01\,M_{\odot}$ would necessitate a sustained MT time of over $5\times 10^{5}$ years. Finally, there is the CE evolution phase. Theoretical studies suggest that the NS could experience hypercritical accretion ($2\times 10^{-4}\,M_{\odot}$ yr-1) in the CE phase due to the loss of accretion energy through neutrino emission (Houck & Chevalier, 1991; Ivanova, 2011). More recent three-dimensional hydrodynamical simulations indicate that the NS can accrete at most a few percent of $1\,M_{\odot}$ due to density gradients in the CE structures (MacLeod & Ramirez-Ruiz, 2015, 2015b). Furthermore, accretion in CE environments typically follows the spherically symmetric Hoyle-Lyttleton model, with no persistent accretion disks forming around the embedded objects (MacLeod et al., 2017; Hutchinson-Smith et al., 2024; Rosselli-Calderon et al., 2024). This aligns with the significant misalignment between the spin and orbital axes observed in PSR J1928+1815 (Yang et al., 2025), suggesting inefficient angular momentum transfer during CE evolution.
@@
In this work we focus on the RLO process proceeding CE evolution, motivated by the observational evidence for super-Eddington accretion onto magnetized NSs in intermediate-mass or high-mass X-ray binaries. Several pulsating ultraluminous X-ray sources (ULXs) have been discovered (Kuranov et al., 2021, for a review). For instance, the ULX in NGC 5907 exhibited a luminosity about 1000 times the Eddington limit (Israel et al., 2017). Other ULXs, such as M82 ULX-2, M51 ULX-8, and NGC 1313 X-2, have been observed with luminosities approaching $\sim 100$ times the Eddington limit (Grisé et al., 2008; Bachetti et al., 2014; Sathyaprakash et al., 2019; Middleton et al., 2019; Brightman et al., 2020). Super-Eddington accretion in NS ULXs is generally attributed to strong NS magnetic fields (e.g., Mushtukov et al., 2015; Dall’Osso et al., 2015) and the accretion disk structure (see below). This implies that, despite the relatively short duration of the RLO phase in NS intermediate-mass or high-mass X-ray binaries, it is still feasible for the NS to accrete a sufficient amount of matter and angular momentum, potentially contributing to its recycling.
@@
In a NS + He star binary system, if the He star is sufficiently massive, it may eventually undergo a supernova (SN) explosion and form a NS. If this process does not disrupt the binary system, a double NS (DNS) system will form. In currently observed DNSs, the first-born NSs have spin periods ranging from 17 to 185 ms (See Deng et al., 2024, Table 1), suggesting that they have undergone substantial recycling. Previous studies proposed that recycling of the first-born NSs generally occurs during the Case BB phase in the NS + He star stage (Tauris et al., 2013a, 2015, 2017; Guo et al., 2025), neglecting accretion before the CE phase. The discovery of PSR J1928+1815 indicates that the latter may also play a role, thus warranting further investigation.
@@ 2 Method and assumptions
To simulate the binary star evolution, we employed both the binary population synthesis code BSE originally developed by Hurley et al. (2002) and updated by Shao & Li (2014) and the MESA stellar evolution code (version 11554 Paxton et al., 2011, 2013, 2015, 2018, 2019). We used the BSE code to model the formation of NS X-ray binaries. We assumed a constant star formation rate of $3\,M_{\odot}$ yr-1 over 12 Gyr. The primordial binaries consisted of a primary star with mass $M_{1}$ and a secondary star with mass $M_{2}$ in circular orbits. The primary star’s mass distribution followed the initial mass function (IMF) of Kroupa et al. (1993), spanning $5-40\,M_{\odot}$. We adopted a uniform distribution between 0 and 1 for the mass ratio ($q=M_{2}/M_{1}$), as suggested by Kobulnicky & Fryer (2007). The initial orbital separation ($a$) was sampled from a logarithmic uniform distribution between 3 and $10^{4}$ $R_{\odot}$, following Abt (1983).
Stellar wind mass loss was modeled using the fitting formula of Nieuwenhuijzen & de Jager (1990). For OB stars with effective temperatures above 11,000 K and for stripped He stars, we used the simulated relations of Vink et al. (2001) and Vink (2017), respectively. We assumed that half of the mass transferred from the primary star via RLO was accreted by the companion star and adopted the corresponding critical mass ratios, $q_{\rm cr}$ calculated by Shao & Li (2014, their Figure 1) as the criterion for stable MT. In our model, if MT becomes dynamically unstable, a CE phase ensues.
To model CE evolution, we compared the binary’s orbital energy to the binding energy of the donor’s envelope (Webbink, 1984)
@@
where $\alpha_{\rm CE}$ represents the CE ejection efficiency, $M_{\rm 1c}$ the mass of the primary’s core, $a_{\rm i}$ and $a_{\rm f}$ the orbital separations of the binary before and after the CE stage respectively, $r_{\rm L}$ the RL radius at the onset of CE, and $\lambda$ the binding energy parameter of the primary’s envelope. If the orbital energy exceeds the binding energy, the envelope is assumed to be ejected. To assess the potential influence of $\alpha_{\rm CE}$, we considered three values: 0.3, 1, and 3. The binding energy parameter, denoted as $\lambda$, which characterizes the structure of the stellar envelope, is influenced by both the mass of the star and its evolutionary stage (Ivanova et al., 2013). Building on the work of Xu & Li (2010b, a), Wang et al. (2016b) used the MESA code to systematically calculate $\lambda$ values, adopting a modified definition of the core-envelope boundary. We extended this approach by using MESA to calculate $\lambda$ values for a denser grid of stellar masses and radii.
@@
For the modeling of NS formation, we incorporated a rapid explosion mechanism to determine remnant masses and natal kicks, following the framework of Fryer et al. (2012). We used the criterion of Fryer et al. (2012) to distinguish between NSs formed in core-collapse supernovae (CCSN) and electron-capture supernovae (ECSN). NSs formed via CCSN were assigned a kick velocity drawn from a Maxwellian distribution with a standard deviation of $\sigma_{\rm CCSN}=$ 265 km s-1 (Hobbs et al., 2005); NSs from ECSN received a lower kick velocity with $\sigma_{\rm ECSN}=$ 20 km s-1 (Podsiadlowski et al., 2004; van den Heuvel, 2004; Verbunt et al., 2017).
@@
Our calculations indicate that in the resulting NS-main-sequence (MS) binary systems, over 60% of the NSs originate from ECSN. Following the prescription of Fryer et al. (2012), the NS mass formed through this channel is uniformly set to 1.38 $M_{\odot}$. Accordingly, in our subsequent detailed evolutionary calculations, we adopted an initial NS mass of 1.4 $M_{\odot}$ for all systems. However, we note that the NS masses generated by the population synthesis simulations exhibit a distribution ranging from $\sim$1.1 to $\sim$2.0 $M_{\odot}$, rather than being entirely concentrated at 1.4 $M_{\odot}$.
@@
For the stellar evolution calculations performed with MESA, each binary system was initialized with a NS and a MS companion of Solar metallicity (Asplund et al., 2009, Z = 0.0142). We adopted the mixing-length theory of Böhm-Vitense (1958), as outlined by Cox & Giuli (1968), to model convection. The mixing-length parameter $\alpha_{\rm MLT}$ was set to 2. Convective regions were identified using the Ledoux criterion (Ledoux, 1947). We incorporated overshooting from the hydrogen-burning convective core using a step-overshooting scheme, extending the core’s size by $\alpha_{\rm ov}$ = 0.335 pressure scale heights in accordance with Brott et al. (2011). For convective cores after the MS, we included a small degree of exponential overshooting (Herwig, 2000) with a decay length scale of $f=0.01$, as overshooting from other convective regions is less well characterized.
@@
We adopted the modified version of the Kolb & Ritter (1990) method to model MT, incorporating the effects of extended atmospheres as described in Ritter (1988) and considering the possibility of optically thick regions overflowing the donor’s RL. The microphysics of accretion imposes several additional constraints on the accretion rate. One such constraint is the Eddington limited accretion. When the accretion luminosity reaches $L_{\rm Edd}=4\pi GM_{\rm NS}c/\kappa$ (where $G$ is the gravitational constant, $M_{\rm NS}$ is NS’s mass, $c$ is the speed of light, and $\kappa$ is the opacity assumed to be due to pure electron scattering), radiation pressure balances gravity, halting steady accretion. This luminosity limit corresponds to an accretion rate limit (Tauris et al., 2013a)
@@
where $X_{\rm H}$ is the H mass fraction of the accreted matter. For super-Eddington accretion, Chashkina et al. (2017) extended the standard thin disk model proposed by Shakura & Sunyaev (1973) to geometrically thick disk accretion. They argued that when the disk’s half-thickness is comparable with the inner radius, the local Eddington limit is reached if the inner radius $R_{0}$ exceeds the spherization radius $R_{\mathrm{sph}}$. Here $R_{0}$ is taken to be the NS magnetospheric radius approximated by the Alfvén radius
@@
where $\dot{M}_{\rm MT}$ denotes the mass transfer rate. This yielded the critical accretion rate (Chashkina et al., 2017)
@@
At higher accretion rates, advection becomes dominant in the disk (Lipunova, 1999), leading to an enhanced critical accretion rate  [^1] The magnitude of $\dot{M}_{\rm crit}$ is contingent upon several additional parameters: the viscosity parameter, the vertical effective polytropic index of the accretion disk, and the energy fraction that powers the disk winds. Here, we adopt the value given by Gao et al. (2022). (Chashkina et al., 2019)
@@
To address the maximum accretion rate ($\dot{M}_{\rm NS,max}$) onto the NS during MT, we followed Chashkina et al. (2017, 2019) and set $\dot{M}_{\rm NS,max}=\dot{M}_{\rm crit}$. If the MT rate exceeds the $\dot{M}_{\rm NS,max}$, we postulated that the excess material is expelled from the binary system in the form of isotropic wind from the NS.
To evaluate the impact of the initial magnetic field $B_{0}$ on the calculated results, we considered three different values: $B_{0}=10^{11}$, $10^{12}$, and $10^{13}$ G, respectively [^2] Due to the strong NS magnetic field, the accreted matter is funneled to the relatively small polar areas of the NS. The temperature and pressure are sufficient for stable thermonuclear burning, precluding the appearance of thermonuclear bursts (Bildsten et al., 1997). Thus, we assumed all the accreted matter remains within the NS.. We adopted an analytical form of magnetic field evolution due to accretion, as given by (Shibazaki et al., 1989; Zhang & Kojima, 2006)
@@
According to Tauris et al. (2012), the relation between the final NS spin period and the accreted mass is
@@
where $I$ is the moment of inertia of the NS, and $K_{\rm su}$ and $K_{\rm sd}$ represent the spin-up and spin-down torques, respectively. We adopted the following prescriptions (Lipunov, 1992; Beskin et al., 1993):
@@
We set $k_{\rm t}=1$ (Shakura et al., 2012) and $\eta=0.8$. In the accretor state, the balance between the spin-up and spin-down torques results in the equilibrium spin period
@@
We set the initial NS mass to $1.4\,M_{\odot}$ and the initial donor mass between 1 and $20\,M_{\odot}$ in steps of 0.5 $M_{\odot}$. The logarithm of the initial binary orbital period was distributed over the range $0.5\leq\log P_{\rm orb,i}({\rm day})\leq 4$ in steps of $\Delta\log P_{\rm orb,i}({\rm day})=0.05$. Following Shao & Li (2021), we assumed that CE evolution occurs if either of the following conditions is satisfied: (1) $\dot{M}_{\rm MT}>\dot{M}_{\rm trap}=2\dot{M}_{\rm Edd}R_{\rm L,NS}/R_{\rm NS}$, where $R_{\rm L,NS}$ is the RL radius of NS (King & Begelman, 1999; Belczynski et al., 2008); and (2) $\dot{M}_{\rm MT}>0.02M_{\rm donor}/P_{\rm orb}$ (Pavlovskii & Ivanova, 2015); and $R_{\rm donor}>R_{\rm L_{2}}$ (Ge et al., 2020). Each binary evolutionary track was terminated if CE evolution occurred or the model numbers exceeded 10,000.

### v2

@@ 1 Introduction
The recent discovery of the pulsar J1928+1815 by the Five-hundred-meter Aperture Spherical radio Telescope (FAST) offers a novel perspective on the formation and evolution of millisecond pulsars (MSPs, Yang2025). This pulsar, with a spin period of 10.55 ms, resides in a binary system with an orbital period of approximately 0.15 day. Its companion is thought to be a helium (He) star with mass $1-1.6\,M_{\odot}$.
@@
Observationally, binary systems containing stripped He stars have been identified in the Magellanic Clouds (Drout2023). In the Milky Way, X-ray binaries with He star donors have also been discovered, with accretors including white dwarfs (WDs, Mereghetti2009) and neutron stars (NSs, Mereghetti2016). However, prior to this discovery, no binary systems containing recycled pulsars and He star companions had been identified. The detection of PSR J1928+1815 fills this observational gap, representing the first confirmed binary system of this kind. It provides crucial evidence for testing and extending existing binary evolution theories and opens a new window for in-depth studies of the accretion-induced spin-up process in MSPs. Through detailed observations and theoretical analyses of this system, we can gain a better understanding of the formation mechanisms and physical properties of MSPs.
@@
MSPs are widely believed to form through the recycling of old NSs via accretion of matter and angular momentum from a companion star in low-mass X-ray binaries (LMXBs, Alpar1982; Radhakrishnan1982). This hypothesis is supported by the detection of millisecond X-ray pulsations in LMXBs and observations of transitions from accreting millisecond X-ray pulsars to MSPs (Wijnands1998; Archibald2009; Papitto2013; Roy2015).
@@
In contrast, systems such as PSR J1928+1815 are thought to originate from intermediate-mass or high-mass X-ray binaries (IMXBs or HMXBs) consisting of a NS and a more massive companion star (e.g., Vigna2018; Shao2018a; Deng2024a). When the companion star evolves to be a giant star and overflows its Roche-lobe, mass transfer (MT) will typically be dynamically unstable because of the large mass ratio. The binary enters the common envelope (CE) phase (Ivanova2013), during which the giant star’s envelope is shed. If the binary survives CE evolution, a recycled NS and a He star, which is the core of the giant star, are left.
@@
Theoretically, accelerating a NS’s spin period to 10 ms requires accretion of at least $0.01\,M_{\odot}$ of matter (e.g., Tauris2012), and a NS in intermediate-mass or high-mass X-ray binaries may undergo multiple accretion phases. Initially, there is accretion from the stellar wind of the giant companion before Roche lobe overflow (RLO). While wind accretion can provide some accreted matter ($\sim 0.005\,M_{\odot}$, Yang2025), its impact on the NS’s spin is uncertain, as it may both accelerate and decelerate the spin (see Tauris2023; Mao2024, and references therein). Current observations of wind-fed X-ray pulsars indicate that the spin periods of NSs range from $\sim 1-10,000$ s (see Wengss2024, Figure 1), with a typical value of $\sim 100$ s (Lyne2012).
@@
Subsequently, there is the RLO phase. During this phase, if the MT rate is constrained by the Eddington limit, accreting $0.01\,M_{\odot}$ would necessitate a sustained MT time of over $5\times 10^{5}$ years. Finally, there is the CE evolution phase. Theoretical studies suggest that the NS could experience hypercritical accretion ($2\times 10^{-4}\,M_{\odot}$ yr-1) in the CE phase due to the loss of accretion energy through neutrino emission (Houck1991; Ivanova2011). More recent three-dimensional hydrodynamical simulations indicate that the NS can accrete at most a few percent of $1\,M_{\odot}$ due to density gradients in the CE structures (MacLeod2015a; MacLeod2015b). Furthermore, accretion in CE environments typically follows the spherically symmetric Hoyle-Lyttleton model, with no persistent accretion disks forming around the embedded objects (MacLeod2017; Hutchinson2024; Rosselli2024). This aligns with the significant misalignment between the spin and orbital axes observed in PSR J1928+1815 (Yang2025), suggesting inefficient angular momentum transfer during CE evolution.
@@
In this work we focus on the RLO process proceeding CE evolution, motivated by the observational evidence for super-Eddington accretion onto magnetized NSs in intermediate-mass or high-mass X-ray binaries. Several pulsating ultraluminous X-ray sources (ULXs) have been discovered (Kuranov2021, for a review). For instance, the ULX in NGC 5907 exhibited a luminosity about 1000 times the Eddington limit (Israel2017Sci). Other ULXs, such as M82 ULX-2, M51 ULX-8, and NGC 1313 X-2, have been observed with luminosities approaching $\sim 100$ times the Eddington limit (Grise2008; Bachetti2014; Sathyaprakash2019; Middleton2019; Brightman2020). Super-Eddington accretion in NS ULXs is generally attributed to strong NS magnetic fields (e.g., Mushtukov2015; DallOsso2015) and the accretion disk structure (see below). This implies that, despite the relatively short duration of the RLO phase in NS intermediate-mass or high-mass X-ray binaries, it is still feasible for the NS to accrete a sufficient amount of matter and angular momentum, potentially contributing to its recycling.
@@
In a NS + He star binary system, if the He star is sufficiently massive, it may eventually undergo a supernova (SN) explosion and form a NS. If this process does not disrupt the binary system, a double NS (DNS) system will form. In currently observed DNSs, the first-born NSs have spin periods ranging from 17 to 185 ms (See Deng2024a, Table 1), suggesting that they have undergone substantial recycling. Previous studies proposed that recycling of the first-born NSs generally occurs during the Case BB phase in the NS + He star stage (Tauris2013; Tauris2015; Tauris2017; Guo2025), neglecting accretion before the CE phase. The discovery of PSR J1928+1815 indicates that the latter may also play a role, thus warranting further investigation.
@@ 2 Method and assumptions
To simulate the binary star evolution, we employed both the binary population synthesis code BSE originally developed by Hurley2002 and updated by Shao2014 and the MESA stellar evolution code (version 11554 Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019). We used the BSE code to model the formation of NS X-ray binaries. We assumed a constant star formation rate of $3\,M_{\odot}$ yr-1 over 12 Gyr. The primordial binaries consisted of a primary star with mass $M_{1}$ and a secondary star with mass $M_{2}$ in circular orbits. The primary star’s mass distribution followed the initial mass function (IMF) of Kroupa1993, spanning $5-40\,M_{\odot}$. We adopted a uniform distribution between 0 and 1 for the mass ratio ($q=M_{2}/M_{1}$), as suggested by Kobulnicky2007. The initial orbital separation ($a$) was sampled from a logarithmic uniform distribution between 3 and $10^{4}$ $R_{\odot}$, following Abt1983.
Stellar wind mass loss was modeled using the fitting formula of Nieuwenhuijzen1990. For OB stars with effective temperatures above 11,000 K and for stripped He stars, we used the simulated relations of Vink2001 and Vink2017, respectively. We assumed that half of the mass transferred from the primary star via RLO was accreted by the companion star and adopted the corresponding critical mass ratios, $q_{\rm cr}$ calculated by Shao2014 as the criterion for stable MT. In our model, if MT becomes dynamically unstable, a CE phase ensues.
To model CE evolution, we compared the binary’s orbital energy to the binding energy of the donor’s envelope (Webbink1984)
@@
where $\alpha_{\rm CE}$ represents the CE ejection efficiency, $M_{\rm 1c}$ the mass of the primary’s core, $a_{\rm i}$ and $a_{\rm f}$ the orbital separations of the binary before and after the CE stage respectively, $r_{\rm L}$ the RL radius at the onset of CE, and $\lambda$ the binding energy parameter of the primary’s envelope. If the orbital energy exceeds the binding energy, the envelope is assumed to be ejected. To assess the potential influence of $\alpha_{\rm CE}$, we considered three values: 0.3, 1, and 3. The binding energy parameter, denoted as $\lambda$, which characterizes the structure of the stellar envelope, is influenced by both the mass of the star and its evolutionary stage (Ivanova2013). Building on the work of Xu2010a; Xu2010b, Wang2016RAA used the MESA code to systematically calculate $\lambda$ values, adopting a modified definition of the core-envelope boundary. We extended this approach by using MESA to calculate $\lambda$ values for a denser grid of stellar masses and radii.
@@
For the modeling of NS formation, we incorporated a rapid explosion mechanism to determine remnant masses and natal kicks, following the framework of Fryer2012. We used the criterion of Fryer2012 to distinguish between NSs formed in core-collapse supernovae (CCSN) and electron-capture supernovae (ECSN). NSs formed via CCSN were assigned a kick velocity drawn from a Maxwellian distribution with a standard deviation of $\sigma_{\rm CCSN}=$ 265 km s-1 (Hobbs2005); NSs from ECSN received a lower kick velocity with $\sigma_{\rm ECSN}=$ 20 km s-1 (Podsiadlowski2004; Heuvel2004; Verbunt2017).
@@
Our calculations indicate that in the resulting NS-main-sequence (MS) binary systems, over 60% of the NSs originate from ECSN. Following the prescription of Fryer2012, the NS mass formed through this channel is uniformly set to 1.38 $M_{\odot}$. Accordingly, in our subsequent detailed evolutionary calculations, we adopted an initial NS mass of 1.4 $M_{\odot}$ for all systems. However, we note that the NS masses generated by the population synthesis simulations exhibit a distribution ranging from $\sim$1.1 to $\sim$2.0 $M_{\odot}$, rather than being entirely concentrated at 1.4 $M_{\odot}$.
@@
For the stellar evolution calculations performed with MESA, each binary system was initialized with a NS and a MS companion of Solar metallicity (Asplund2009, Z = 0.0142). We adopted the mixing-length theory of Bohm1958, as outlined by Cox1968, to model convection. The mixing-length parameter $\alpha_{\rm MLT}$ was set to 2. Convective regions were identified using the Ledoux criterion (Ledoux1947). We incorporated overshooting from the hydrogen-burning convective core using a step-overshooting scheme, extending the core’s size by $\alpha_{\rm ov}$ = 0.335 pressure scale heights in accordance with Brott2011. For convective cores after the MS, we included a small degree of exponential overshooting (Herwig2000) with a decay length scale of $f=0.01$, as overshooting from other convective regions is less well characterized.
@@
We adopted the modified version of the Kolb1990 method to model MT, incorporating the effects of extended atmospheres as described in Ritter1988 and considering the possibility of optically thick regions overflowing the donor’s RL. The microphysics of accretion imposes several additional constraints on the accretion rate. One such constraint is the Eddington limited accretion. When the accretion luminosity reaches $L_{\rm Edd}=4\pi GM_{\rm NS}c/\kappa$ (where $G$ is the gravitational constant, $M_{\rm NS}$ is NS’s mass, $c$ is the speed of light, and $\kappa$ is the opacity assumed to be due to pure electron scattering), radiation pressure balances gravity, halting steady accretion. This luminosity limit corresponds to an accretion rate limit (Tauris2013)
@@
where $X_{\rm H}$ is the H mass fraction of the accreted matter. For super-Eddington accretion, Chashkina2017 extended the standard thin disk model proposed by Shakura1973 to geometrically thick disk accretion. They argued that when the disk’s half-thickness is comparable with the inner radius, the local Eddington limit is reached if the inner radius $R_{0}$ exceeds the spherization radius $R_{\mathrm{sph}}$. Here $R_{0}$ is taken to be the NS magnetospheric radius approximated by the Alfvén radius
@@
where $\dot{M}_{\rm MT}$ denotes the mass transfer rate. This yielded the critical accretion rate (Chashkina2017)
@@
At higher accretion rates, advection becomes dominant in the disk (Lipunova1999), leading to an enhanced critical accretion rate  [^1] The magnitude of $\dot{M}_{\rm crit}$ is contingent upon several additional parameters: the viscosity parameter, the vertical effective polytropic index of the accretion disk, and the energy fraction that powers the disk winds. Here, we adopt the value given by Gao2022. (Chashkina2019)
@@
To address the maximum accretion rate ($\dot{M}_{\rm NS,max}$) onto the NS during MT, we followed Chashkina2017; Chashkina2019 and set $\dot{M}_{\rm NS,max}=\dot{M}_{\rm crit}$. If the MT rate exceeds the $\dot{M}_{\rm NS,max}$, we postulated that the excess material is expelled from the binary system in the form of isotropic wind from the NS.
To evaluate the impact of the initial magnetic field $B_{0}$ on the calculated results, we considered three different values: $B_{0}=10^{11}$, $10^{12}$, and $10^{13}$ G, respectively [^2] Due to the strong NS magnetic field, the accreted matter is funneled to the relatively small polar areas of the NS. The temperature and pressure are sufficient for stable thermonuclear burning, precluding the appearance of thermonuclear bursts (Bildsten1997). Thus, we assumed all the accreted matter remains within the NS.. We adopted an analytical form of magnetic field evolution due to accretion, as given by (Shibazaki1989; Zhang2006)
@@
According to Tauris2012, the relation between the final NS spin period and the accreted mass is
@@
where $I$ is the moment of inertia of the NS, and $K_{\rm su}$ and $K_{\rm sd}$ represent the spin-up and spin-down torques, respectively. We adopted the following prescriptions (Lipunov1992; Beskin1993):
@@
We set $k_{\rm t}=1$ (Shakura2012) and $\eta=0.8$. In the accretor state, the balance between the spin-up and spin-down torques results in the equilibrium spin period
@@
We set the initial NS mass to $1.4\,M_{\odot}$ and the initial donor mass between 1 and $20\,M_{\odot}$ in steps of 0.5 $M_{\odot}$. The logarithm of the initial binary orbital period was distributed over the range $0.5\leq\log P_{\rm orb,i}({\rm day})\leq 4$ in steps of $\Delta\log P_{\rm orb,i}({\rm day})=0.05$. Following Shao2021, we assumed that CE evolution occurs if either of the following conditions is satisfied: (1) $\dot{M}_{\rm MT}>\dot{M}_{\rm trap}=2\dot{M}_{\rm Edd}R_{\rm L,NS}/R_{\rm NS}$, where $R_{\rm L,NS}$ is the RL radius of NS (King1999; Belczynski2008); and (2) $\dot{M}_{\rm MT}>0.02M_{\rm donor}/P_{\rm orb}$ (Pavlovskii2015); and $R_{\rm donor}>R_{\rm L_{2}}$ (Ge2020). Each binary evolutionary track was terminated if CE evolution occurred or the model numbers exceeded 10,000.